Caspian Sea Drinks is considering the purchase of a new water filtration system produced by Rube Goldberg Machines. This new equipment, the RGM-7000, will allow Caspian Sea Drinks to expand production. It will cost $13.00 million fully installed and will be fully depreciated over a 20 year life, then removed for no cost. The RGM-7000 will result in additional revenues of $3.74 million per year and increased operating costs of $568,841.00 per year. Caspian Sea Drinks' marginal tax rate is 23.00%. If Caspian Sea Drinks uses a 12.00% discount rate, then the net present value of the RGM-7000 is _____.
Answer #1

Annual depreciation = 13,000,000 / 20 = 650,000

OCF from year 1 to year 20 = (Revenue - costs - depreciation)(1 - tax) + depreciation

OCF from year 1 to year 20 = (3,740,000 - 568,841 - 650,000)(1 - 0.23) + 650,000

OCF from year 1 to year 20 = 1,941,292.43 + 650,000

OCF from year 1 to year 20 = $2,591,292.43

Net present value = Present value of cash inflows - present value of cash outflows

Net present value = Annuity * [1 - 1 / (1 + r)n] / r - Initial investment

Net present value = 2,591,292.43 * [1 - 1 / (1 + 0.12)20] / 0.12 - 13,000,000

Net present value = 2,591,292.43 * [1 - 0.103667] / 0.12 - 13,000,000

Net present value = 2,591,292.43 * 7.469444- 13,000,000

Net present value = $6,355,512.72